find two consecutive integers such that the sum of their squares is 61.
Answer by checkley71(8403)

THE NUMBERS ARE X AND X+1
THE FORMULA IS X~2+(X+1)~2=61 OR X~2+X~2+2X+1=61 OR 2X~2+2X-60=0 OR
X~2+X-60=0 OR (X+6)(X-5)=0 OR X=-6, X=+5 AND X+1=6
PROOF 5~2+6~2=61 OR 25+36=61 OR 61=61